# Shuangrui Ding1  Rui Qian1  Xiaoyi Dong1,2  Pan Zhang2 Yuhang Zang2  Yuhang Cao2  Yuwei Guo1   Dahua Lin1   Jiaqi Wang2

### v2

@@ 3.3 Object-aware Memory Bank Construction
Finally, the modulated memory keys ~$\widetilde{\mathcal{M}}_{t}=\{\widetilde{\mathbf{M}}_{\tau}\}_{\tau\in\mathcal{I}}$ are used in the memory module’s cross-attention mechanism to update the current frame’s features. By using the available occlusion scores as indicators, we effectively emphasize memory entries with more reliable object cues while introducing minimal computational overhead.
@@ 4.1 Datasets
To evaluate our method, we select 6 standard VOS benchmarks and report the following metrics: $\mathcal{J}$ (region similarity), $\mathcal{F}$ (contour accuracy), and the combined $\mathcal{J}\&\mathcal{F}$. All evaluations are conducted in a semi-supervised setting, where the first-frame mask is provided. The datasets used for testing are detailed as follows:
SA-V (Ravi et al., 2024) is a large-scale video segmentation dataset designed for promptable visual segmentation across diverse scenarios. It encompasses 50.9K video clips, aggregating to 642.6K masklets with 35.5M meticulously annotated masks. The dataset presents a challenge with its inclusion of small, occluded, and reappearing objects throughout the videos. The dataset is divided into training, validation, and testing sets, with most videos allocated to the training set for robust model training. The validation set has 293 masklets across 155 videos for model tuning, while the testing set includes 278 masklets across 150 videos for comprehensive evaluation.
LVOS v1 (Hong et al., 2023) is a VOS benchmark for long-term video object segmentation in realistic scenarios. It comprises 720 video clips with 296,401 frames and 407,945 annotations, with an average video duration of over 60 seconds. LVOS introduces challenging elements such as long-term object reappearance and cross-temporal similar objects. In LVOS v1, the dataset includes 120 videos for training, 50 for validation, and 50 for testing.
LVOS v2 (Hong et al., 2024) expends LVOS v1 and provides 420 videos for training, 140 for validation, and 160 for testing. This paper primarily utilizes v2, as it already includes the sequences present in v1. The dataset spans 44 categories, capturing typical everyday scenarios, with 12 of these categories deliberately left unseen to evaluate and better assess the generalization capabilities of VOS models.
MOSE (Ding et al., 2023a) is a challenging VOS dataset targeted on complex, real-world scenarios, featuring 2,149 video clips with 431,725 high-quality segmentation masks. These videos are split into 1,507 training videos, 311 validation videos, and 331 testing videos.
VOST (Tokmakov et al., 2023) is a semi-supervised video object segmentation benchmark that emphasizes complex object transformations. Unlike other datasets, VOST includes objects that are broken, torn, or reshaped, significantly altering their appearance. It comprises more than 700 high-resolution videos, captured in diverse settings, with an average duration of 21 seconds, all densely labeled with instance masks.
PUMaVOS (Bekuzarov et al., 2023) is a novel video dataset designed for benchmarking challenging segmentation tasks. It includes 24 video clips, each ranging from 13.5 to 60 seconds (28.7 seconds on average) at 480p resolution with varying aspect ratios. PUMaVOS focuses on difficult scenarios where annotation boundaries do not align with clear visual cues, such as half faces, necks, tattoos, and pimples, commonly encountered in video production.
4.2 Main Results
SAM2Long consistently improves SAM 2 over all model sizes and datasets. Table 1 presents an overall comparison between SAM 2 and SAM2Long across various model sizes on the SA-V validation and test sets, as well as the LVOS v2 validation set. In total, the table includes 8 model variations, covering SAM 2 and the latest SAM 2.1 across four model sizes. The average performance across 24 experiments shows an improvement of 3 points in the $\mathcal{J}\&\mathcal{F}$ score. These results confirm that SAM2Long consistently outperforms the SAM 2 baseline by a large margin. For instance, for SAM2Long-Large, achieves an improvement of 4.5 and 5.3 over SAM 2 on SA-V val and test sets. This pattern is also evident in the LVOS validation set, where SAM2Long delivers notable performance gains over SAM 2 for each model size. These results showcase the effectiveness of the training-free memory tree in long-time video segmentation scenarios.
SAM2Long outperforms previous methods and excels in unseen categories. We also compare our proposed method, SAM2Long, with various state-of-the-art VOS methods on both the SA-V (Ravi et al., 2024) and LVOS (Hong et al., 2023; 2024) datasets, as shown in Table 2 and 3. Although SAM 2.1 already surpasses previous methods by a large margin, SAM2.1Long pushes these limits even further. Specifically, our method achieves a $\mathcal{J}\&\mathcal{F}$ score of 81.1 on the SA-V validation set, a 2.5-point improvement over SAM 2.1. For LVOS, SAM2.1Long attains $\mathcal{J}\&\mathcal{F}$ scores of 83.4 and 85.9 on the v1 and v2 subsets, respectively, outperforming SAM 2.1 by 3.2 and 1.8 points. Notably, SAM2Long particularly excels in unseen categories, achieving $\mathcal{J}$ and $\mathcal{F}$ scores of 79.1 and 86.2. The significant improvements of 7.5 and 5.1 points over SAM 2 highlight its robust generalization capabilities.
SAM2Long demonstrates versatility when handling videos with various challenges. In addition to the SA-V and LVOS datasets, we evaluate our proposed SAM2Long on other VOS benchmarks in Table 4. On the PUMaVOS (Bekuzarov et al., 2023), which challenges models with difficult visual cues, SAM2.1Long outperforms SAM 2.1 with a score of 82.4 compared to 81.1, demonstrating its enhanced ability to handle subtle and ambiguous segmentation tasks. Similarly, on the VOST (Tokmakov et al., 2023), which focuses on objects undergoing extreme transformations, SAM2.1Long shows improvement with a $\mathcal{J}\&\mathcal{F}$ score of 54.0, a nearly 1-point gain over SAM2.1. On the MOSE (Ding et al., 2023a), which involves complex real-world scenes, SAM2.1Long achieves a $\mathcal{J}\&\mathcal{F}$ score of 75.2, surpassing SAM 2.1’s score of 74.5. Given that the stronger SAM 2.1-L shows no improvement over the SAM 2-L model on the MOSE benchmark, our performance gain with SAM2.1Long is particularly notable. However, YouTube-VOS Xu et al. (2018), with an average duration of only 4.5 seconds, offers limited opportunities to showcase long-term tracking capabilities. Here, both SAM2.1 and SAM2.1Long achieve the same score of 88.7. These results underscore that we have preserved the fundamental segmentation capabilities of SAM 2 while enhancing its long-term abilities, demonstrating the robustness and versatility of SAM2Long across a range of VOS benchmarks.
Table 1: Performance comparison on SA-V (Ravi et al., 2024) and LVOS v2 (Hong et al., 2024) datasets between SAM 2 and SAM2Long across all model sizes. $\dagger$ We report the re-produced performance of SAM 2 using its open-source code and checkpoint.
Method SA-V val SA-V test LVOS v2 val
[…]
SAM2-T 73.5 70.1 76.9 74.6 71.1 78.0 77.8 74.5 81.2
SAM2.1-T 75.1 71.6 78.6 76.3 72.7 79.8 81.6 77.9 85.2
[…]
SAM2.1-S 76.9 73.5 80.3 76.9 73.3 80.5 82.1 78.6 85.6
[…]
SAM2.1-B+ 78.0 74.6 81.5 77.7 74.2 81.2 83.1 79.6 86.5
[…]
SAM2Long-L 80.8 (4.5$\uparrow$) 77.1 84.5 80.8 (5.3$\uparrow$) 76.8 84.7 85.2 (2.2$\uparrow$) 81.8 88.7
SAM2.1-L 78.6 75.1 82.0 79.6 76.1 83.2 84.0 80.7 87.4
@@ 4.2 Main Results
Method SA-V val SA-V test
[…]
STCN (Cheng et al., 2021) 61.0 57.4 64.5 62.5 59.0 66.0
RDE (Li et al., 2022) 51.8 48.4 55.2 53.9 50.5 57.3
SwinB-AOT (Yang et al., 2021a) 51.1 46.4 55.7 50.3 46.0 54.6
SwinB-DeAOT (Yang & Yang, 2022) 61.4 56.6 66.2 61.8 57.2 66.3
XMem (Cheng & Schwing, 2022) 60.1 56.3 63.9 62.3 58.9 65.8
DEVA (Cheng et al., 2023) 55.4 51.5 59.2 56.2 52.4 60.1
Cutie-base+ Cheng et al. (2024) 61.3 58.3 64.4 62.8 59.8 65.8
[…]
SAM 2.1 (Ravi et al., 2024) 78.6 75.1 82.0 79.6 76.1 83.2
SAM2Long (ours) 79.7 74.7 84.7 80.8 76.8 84.7
SAM2.1Long (ours) 81.1 77.5 84.7 81.2 77.6 84.9
@@
LWL (Bhat et al., 2020) 56.4 51.8 60.9 60.6 58.0 64.3 57.2 62.9
CFBI (Yang et al., 2020) 51.5 46.2 56.7 55.0 52.9 59.2 51.7 56.2
STCN (Cheng et al., 2021) 48.9 43.9 54.0 60.6 57.2 64.0 57.5 63.8
RDE (Li et al., 2022) 53.7 48.3 59.2 62.2 56.7 64.1 60.8 67.2
DeAOT (Yang et al., 2021a) - - - 63.9 61.5 69.0 58.4 66.6
XMem (Cheng & Schwing, 2022) 52.9 48.1 57.7 64.5 62.6 69.1 60.6 65.6
DDMemory (Hong et al., 2023) 60.7 55.0 66.3 - - - - -
@@
SAM 2.1 (Ravi et al., 2024) 80.2 75.4 84.9 84.1 80.7 87.4 80.6 87.7
SAM2Long (ours) 81.3 76.4 86.2 84.2 82.3 89.2 79.1 86.2
SAM2.1Long (ours) 83.4 78.4 88.5 85.9 81.7 88.6 83.0 90.5
Dataset Duration SAM 2.1 SAM2.1Long
seconds $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$
PUMaVOS (Bekuzarov et al., 2023) 28.7 81.1 82.4
VOST (Tokmakov et al., 2023) 21 53.1 54.0
MOSE (Ding et al., 2023a) 12.4 74.5 75.2
YTVOS (Xu et al., 2018) 4.5 88.7 88.7
@@ 4.3 Ablation Study
We conduct a series of ablation studies on the validation split of SA-V dataset and use SAM2-Large as default model size.
Number of Memory Pathways $P$. We ablate the number of memory pathways to assess their impact on SAM2Long in Table 8. Note that setting $P=1$ reverts to the SAM 2 baseline. Increasing the number of memory pathways to $P=2$ yields a notable improvement, raising the $\mathcal{J}\&\mathcal{F}$ score to 80.1. This result demonstrates that the proposed memory tree effectively boosts the model’s ability to track the correct object while reducing the impact of occlusion. Further increasing the number of memory pathways to $P=3$ achieves the best performance. However, using $P=4$ shows no additional gains, suggesting that three pathways strike the optimal balance between accuracy and computational efficiency for the SAM 2 model.
In terms of speed, since we maintain a fixed number of memory pathways at every time step, the processing speed remains efficient. Using three memory pathways results in only an $18\%$ slowdown and an $8\%$ increase in GFlops, while delivering a 4.5-point improvement in performance.
Iou Threshold $\delta_{\text{iou}}$. The choice of the IoU threshold $\delta_{\text{iou}}$ is crucial for selecting frames with reliable object cues. As shown in Table 8, setting $\delta_{\text{iou}}=0.3$ yields the highest $\mathcal{J}\&\mathcal{F}$, indicating an effective trade-off between filtering out poor-quality frames and retaining valuable segmentation information. In contrast, having no requirement on mask quality and feeding all masks containing objects into memory ($\delta_{\text{iou}}=0$) decreases the score to 80.0, showing that unreliable frames with poor segmentation harm the SAM 2 model. Meanwhile, an overly strict selection ($\delta_{\text{iou}}=0.9$) degrades performance even more severely to 77.8, as it excludes too many potentially important neighboring frames, causing the model to rely on frames that are too far away from the current frame as memory.
Uncertainty Threshold $\delta_{\text{conf}}$. The uncertainty threshold $\delta_{\text{conf}}$ controls the selection of hypotheses under uncertain conditions. Our results in Table 8 indicate that setting $\delta_{\text{conf}}$ to 2 provides the highest $\mathcal{J}\&\mathcal{F}$ score, indicating an optimal level for uncertainty handling. Lower values (e.g., 0.5) result in suboptimal performance, as they may prematurely commit to incorrect segmentation hypotheses, leading to significant performance drops due to error propagation. On the other hand, higher values (e.g., 5) do not further improve performance, suggesting that beyond a certain threshold, the model does not benefit from additional mask diversity and can efficiently rely on the top-scoring masks when the segmentation is confident.
Memory Attention Modulation $[w_{\text{low}},w_{\text{high}}]$. We explore the effect of modulating the attention weights for memory entries using different ranges in Table 8. The configuration $\left[1,1\right]$ means no modulation is applied. We find that the configuration of $\left[0.95,1.05\right]$ achieves the best performance while increasing the modulation range to ($\left[0.9,1.1\right]$) slightly decreases performance. This result indicates that slight modulation sufficiently emphasizes reliable memory entries.
Table 5: Ablation on number of pathways $P$.
@@
1 76.3 73.0 79.5 1$\times$ 844.1
2 80.1 76.7 83.5 0.93$\times$ 878.2
3 80.8 77.1 84.5 0.82$\times$ 912.3
4 80.7 77.0 84.5 0.75$\times$ 946.4

### v3

@@ 3.3 Object-aware Memory Bank Construction
Finally, the modulated memory keys ~$\widetilde{\mathcal{M}}_{t}=\{\widetilde{\mathbf{M}}_{\tau}\}_{\tau\in\mathcal{I}}$ are used in the memory module’s cross-attention mechanism to update the current frame’s features. By using the available occlusion scores as indicators, we effectively emphasize memory entries with more reliable object cues.
@@ 4.1 Experiments Setup
Datasets. To evaluate our method, we select nine standard VOS benchmarks: SA-V val and test [50], LVOS v1 [26], LVOS v2 [27], MOSE [17], VOST [54], PUMaVOS [4], DAVIS [47], and YTVOS [63]. We report the following metrics: $\mathcal{J}$ (region similarity), $\mathcal{F}$ (contour accuracy), and the combined $&\mathcal{J}\&\mathcal{F}$. All evaluations are conducted in a semi-supervised setting, where the first-frame mask is provided. Details on the benchmarks are provided in the Appendix.
Evaluation. SAM2Long is modified from the official codebase of SAM2 and we report the re-produced performance of SAM2 with the same official code and checkpoint. We noticed a slight difference between the re-produced results and the performance reported in the SAM2 paper, but for a strictly fair comparison, we tend to report both the SAM2 and SAM2Long results under the same hardware/software environment and settings. We set the uncertainty threshold to $\delta_{\text{conf}}=2$, IoU threshold to $\delta_{\text{IoU}}=0.3$, modulation weight to $\left[w_{\text{low}},w_{\text{high}}\right]=[0.95,1.05]$. We find that the performance remains fairly robust across datasets when tuning these hyperparameters; thus, we use the same values to report results for all experiments. The detailed findings are provided in the Appendix for reference.
   Method    SA-V val    SA-V test
   $&\mathcal{J}\&\mathcal{F}$    $\mathcal{J}$    $\mathcal{F}$    $&\mathcal{J}\&\mathcal{F}$    $\mathcal{J}$    $\mathcal{F}$
   SwinB-DeAOT [65]    61.4    56.6    66.2    61.8    57.2    66.3
   DEVA [13]    55.4    51.5    59.2    56.2    52.4    60.1
   Cutie-base+ [14]    61.3    58.3    64.4    62.8    59.8    65.8
   STCN [12]    61.0    57.4    64.5    62.5    59.0    66.0
   XMem [11]    60.1    56.3    63.9    62.3    58.9    65.8
   SAM 2 [50]    76.1    72.9    79.2    76.0    72.6    79.3
   SAM2Long (ours)    80.8    74.7    84.7    80.8    76.8    84.7
   SAM 2.1 [50]    78.6    75.1    82.0    79.6    76.1    83.2
   SAM2.1Long (ours)    81.1    77.5    84.7    81.2    77.6    84.9
@@
LWL [5] 56.4 51.8 60.9 60.6 58.0 64.3 57.2 62.9
RDE [34] 53.7 48.3 59.2 62.2 56.7 64.1 60.8 67.2
STCN [12] 48.9 43.9 54.0 60.6 57.2 64.0 57.5 63.8
XMem [11] 52.9 48.1 57.7 64.5 62.6 69.1 60.6 65.6
DDMemory [26] 60.7 55.0 66.3 - - - - -
SAM 2 [50] 77.9 73.1 82.7 79.8 80.0 86.6 71.6 81.1
SAM2Long (ours) 81.3 76.4 86.2 84.2 82.3 89.2 79.1 86.2
SAM 2.1 [50] 80.2 75.4 84.9 84.1 80.7 87.4 80.6 87.7
SAM2.1Long (ours) 83.4 78.4 88.5 85.9 81.7 88.6 83.0 90.5
4.2 Main Results
SAM2Long consistently improves SAM 2 over all model sizes and datasets. Table 1 presents an overall comparison between SAM 2 and SAM2Long across various model sizes on the SA-V validation and test sets, as well as the LVOS v2 validation set. In total, the table includes 8 model variations, covering SAM 2 and the latest SAM 2.1 across four model sizes. The average performance across 12 experiments shows an improvement of 3.7 points in the $&\mathcal{J}\&\mathcal{F}$ score. These results confirm that SAM2Long consistently outperforms the SAM 2 baseline by a large margin. For instance, for SAM2Long-Large, achieves an improvement of 4.5 and 5.3 over SAM 2 on SA-V val and test sets.
SAM2Long outperforms previous methods. We also compare our proposed method, SAM2Long, with various state-of-the-art VOS methods on both the SA-V [50] and LVOS [26, 27] datasets, as shown in Table 2 and 3. Previous methods such as STCN[12] and XMem[11] adopt memory-based strategies to improve performance and efficiency. STCN models object relations using a single affinity matrix, whereas XMem maintains a three-tier hierarchical memory structure to balance accuracy and GPU usage. While both methods focus on learning feature-based memory representations, SAM2Long takes a different approach by refining memory at the mask level rather than the feature level. Our method efficiently selects the optimal sequence of memory masks using the proposed tree structure, eliminating the need to retrain the original memory features. As a result, SAM2Long surpasses XMem and STCN by over 20 points on both LVOS v2 and SA-V val, demonstrating its superior ability to handle long-term object segmentation.
Dataset Duration SAM 2.1 SAM2.1Long
seconds $&\mathcal{J}\&\mathcal{F}$ $&\mathcal{J}\&\mathcal{F}$
LVOS v1 [26] 95.4 80.2 83.4 (3.2$\uparrow$)
LVOS v2 [27] 68.4 84.1 85.9 (1.8$\uparrow$)
SA-V val [50] 13.8 78.6 81.1 (2.5$\uparrow$)
SA-V test [50] 13.8 79.6 81.2 (1.6$\uparrow$)
PUMaVOS [4] 28.7 81.1 82.4 (1.2$\uparrow$)
VOST [54] 21 53.1 54.0 (0.9$\uparrow$)
MOSE [17] 12.4 74.5 75.2 (0.7$\uparrow$)
YTVOS [63] 4.5 88.7 88.8 (0.1$\uparrow$)
DAVIS-17 [47] 1.8 90.1 90.2 (0.1$\uparrow$)
Performance gains of SAM2Long positively correlate with video duration. The results shown in the Table 4 reveal a clear trend: SAM2Long tends to exhibit greater performance gains with longer videos. Although SAM2Long performs comparably to SAM 2 on shorter-duration datasets [63, 47], this aligns with the limited long-term tracking opportunities provided by short video sequences. However, SAM2Long can still maintain the strong segmentation performance of SAM 2.
4.3 Additional Results on VOT Benchmarks
To further evaluate the effectiveness of SAM2Long, we also apply it to three widely used single-object tracking benchmarks, namely LaSOT [22], LaSOText [23], and GOT-10k [30]. Without whistles and bells, we first generate an initial mask from the bounding box provided for the first frame using the SAM2 Mask decoder. Subsequently, all subsequent masks are transformed into bounding box format by identifying the minimum and maximum coordinates of their non-zero indices. SAM2Long achieves competitive performance with state-of-the-art VOT methods and SAM 2 baseline, demonstrating its strong generalization ability across various tracking scenarios.
@@ 4.3 Additional Results on VOT Benchmarks
KeepTrack [41] 67.1 48.2 -
TOMP [42] 68.5 - -
DropTrack [60] 71.8 52.7 75.9
SeqTrack [9] 72.5 50.7 74.8
MixFormer [16] 70.1 - 71.2
GRM-256 [24] 69.9 - 73.4
ROMTrack [8] 71.4 51.3 74.2
OSTrack [69] 71.1 50.5 73.7
DiffusionTrack [39] 72.3 - 74.7
ODTrack [72] 74.0 53.9 78.2
LORAT [38] 75.1 56.6 78.2
SAM 2 [50] 70.0 56.9 80.7
SAM2Long (ours) 73.9 60.9 81.1
Table 5: Performance comparison with state-of-the-art VOT methods on LaSoT, LaSoText, and GOT-10k benchmarks. The results are reported in terms of AUC (Area Under the success rate Curve) for LaSoT and LaSoText, and AO (Average Overlap) for GOT-10k.
Figure 3: Qualitative comparison between SAM 2 and SAM2Long, with GT (Ground Truth) provided for reference. The last row shows a failure case. Best viewed when zoomed in.
@@ 4.4 Ablation Study
We perform a series of ablation studies on the validation subsets of the SA-V and LVOS v2 datasets, utilizing SAM2-Large as the default model size.
Number of Memory Pathways $P$. We ablate the number of memory pathways to assess their impact on SAM2Long in Table 6. Note that setting $P=1$ reverts to the SAM 2 baseline. Increasing the number of memory pathways to $P=2$ yields a notable improvement, raising the $&\mathcal{J}\&\mathcal{F}$ score to 80.1. This result demonstrates that the proposed memory tree effectively boosts the model’s ability to track the correct object while reducing the impact of occlusion. Further increasing the number of memory pathways to $P=3$ achieves the best performance. However, using $P=4$ shows no additional gains, suggesting that three pathways strike the optimal balance between accuracy and computational efficiency for the SAM 2 model.
@@
$P$ SA-V LVOS FPS GFlops Memory
1 76.3 83.0 22 844.1 5.1GB
2 80.1 85.0 21 878.2 5.2GB
3 80.8 85.4 19 912.3 5.3GB
4 80.7 85.2 17 946.4 5.4GB
Table 6: Ablation study on number of pathways $P$. We report $&\mathcal{J}\&\mathcal{F}$ performance along with FPS, computational cost (GFlops), and memory usage. Throughput and memory consumption are measured on a RTX 3090 GPU with 24 GB of memory.
Memory Frame Selection. Building on IoU-based memory frame selection, we incorporate two additional mechanisms: temporal and spatial selection. The temporal method applies a time decay factor to the IoU score, prioritizing frames closer in time to the current frame. The spatial method integrates frame-wise feature similarity into the IoU score, favoring frames with similar spatial contexts. However, as shown in Table 7, neither approach enhances performance, showing that IoU filtering remains the key factor.
Memory Selection SA-V LVOS
Only by IoU 80.8 85.4
w. temporal 80.4 85.2
w. spatial 80.7 84.8
Table 7: Ablation study on memory frame selection methods. We compare the standard IoU-based selection with additional temporal and spatial selection mechanisms.
Rounding Predicted IoU on Mask Diversity. To evaluate the impact of rounding predicted IoU in selecting diverse masks, we compute the actual IoU between candidate masks chosen based on rounded predicted IoU. As shown in Table 8, rounding to two decimal places significantly reduces the actual IoU, highlighting increased variation among candidates. This enhanced diversity benefits our tree search strategy and improves performance in handling uncertain cases.
Rounding SA-V LVOS actual pairwise IoU
80.8 85.4 51.4
80.4 84.4 84.5
Table 8: Ablation study on rounding predicted IoU.
4.5 Qualitative Comparison
We present a qualitative comparison between SAM 2 and SAM2Long in Figure 3. As illustrated in Figure 3, SAM 2 successfully tracks the correct person in a group of dancing people. However, when occlusion occurs, SAM 2 mistakenly switches to tracking a different, misleading individual. In contrast, SAM2Long handles this ambiguity effectively. Even during the occlusion, SAM2Long manages to resist the tracking error and correctly resumes tracking the original dancer when she reappears. We also present a failure case. When the video features dynamic background changes and distracting elements, SAM2Long struggles to maintain accurate tracking. The model either mistakenly tracks the wrong shirt or completely loses the target. We attribute this to SAM 2’s over-reliance on fine-grained visual details and its lack of semantic understanding.
5 Conclusion
In this paper, we introduce SAM2Long, a training-free enhancement to SAM 2 that alleviates its limitations in long-term video object segmentation. By employing a constrained tree memory structure with object-aware memory modulation, SAM2Long effectively mitigates error accumulation and improves robustness against occlusions, resulting in a more reliable segmentation process over extended periods. Extensive evaluations on six VOS benchmarks demonstrate that SAM2Long consistently outperforms SAM 2, especially in complex video scenarios. Notably, SAM2Long achieves up to a 5-point improvement in $&\mathcal{J}\&\mathcal{F}$ scores on challenging long-term video benchmarks such SA-V and LVOS.
Limitations. While SAM2Long is effective, it has a few limitations. First, its performance is constrained by the inherent capacity of SAM2, as no learnable parameters are modified, restricting its optimization potential. Second, the method primarily focuses on single-object tracking rather than carefully designed multi-object scenarios. Researching specific multi-object techniques will be a promising future direction. However, our experiments demonstrate that SAM2Long already performs exceptionally well in both single-object and multi-object cases. We list the comparison in the Appendix.
Acknowledgemnts
This work was supported by National Key R&D Program of China 2022ZD0161600, Shanghai Artificial Intelligence Laboratory, Hong Kong RGC TRS T41-603/20-R, the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of CPII under the InnoHK.
Appendix A Dataset Information
SA-V [50] is a large-scale video segmentation dataset designed for promptable visual segmentation across diverse scenarios. It encompasses 50.9K video clips, aggregating to 642.6K masklets with 35.5M meticulously annotated masks. The dataset presents a challenge with its inclusion of small, occluded, and reappearing objects throughout the videos. The dataset is divided into training, validation, and testing sets, with most videos allocated to the training set for robust model training. The validation set has 293 masklets across 155 videos for model tuning, while the testing set includes 278 masklets across 150 videos for comprehensive evaluation.
LVOS v1 [26] is a VOS benchmark for long-term video object segmentation in realistic scenarios. It comprises 720 video clips with 296,401 frames and 407,945 annotations, with an average video duration of over 60 seconds. LVOS introduces challenging elements such as long-term object reappearance and cross-temporal similar objects. In LVOS v1, the dataset includes 120 videos for training, 50 for validation, and 50 for testing.
LVOS v2 [27] expends LVOS v1 and provides 420 videos for training, 140 for validation, and 160 for testing. This paper primarily utilizes v2, as it already includes the sequences present in v1. The dataset spans 44 categories, capturing typical everyday scenarios, with 12 of these categories deliberately left unseen to evaluate and better assess the generalization capabilities of VOS models.
MOSE [17] is a challenging VOS dataset targeted on complex, real-world scenarios, featuring 2,149 video clips with 431,725 high-quality segmentation masks. These videos are split into 1,507 training videos, 311 validation videos, and 331 testing videos.
VOST [54] is a semi-supervised video object segmentation benchmark that emphasizes complex object transformations. Unlike other datasets, VOST includes objects that are broken, torn, or reshaped, significantly altering their appearance. It comprises more than 700 high-resolution videos, captured in diverse settings, with an average duration of 21 seconds, all densely labeled with instance masks.
PUMaVOS [4] is a novel video dataset designed for benchmarking challenging segmentation tasks. It includes 24 video clips, each ranging from 13.5 to 60 seconds (28.7 seconds on average) at 480p resolution with varying aspect ratios. PUMaVOS focuses on difficult scenarios where annotation boundaries do not align with clear visual cues, such as half faces, necks, tattoos, and pimples, commonly encountered in video production.
YouTubeVOS-2019 [63] is a large-scale video object segmentation dataset featuring 3,252 sequences with detailed annotations at 6 FPS across 78 diverse categories, including humans, animals, vehicles, and accessories. Each video clip is between 3 to 6 seconds long and frequently contains multiple objects, which have been manually segmented by professional annotators.
DAVIS2017 [47] is a well-known benchmark dataset comprising 60 training videos and 30 validation videos, with a total of 6,298 frames. It offers high-quality, pixel-level annotations for every frame, making it a standard resource for evaluating different VOS methods.
LaSOT [22] is a large-scale tracking dataset designed for long-term visual object tracking. It contains 1,400 videos spanning 70 object categories, with an average sequence length exceeding 2,500 frames, making it a challenging benchmark for evaluating tracking algorithms. LaSOText[23] is an extension of LaSOT, featuring a subset of 15 categories with 150 videos.
GOT-10k [30] is a large-scale generic object tracking dataset that covers over 10,000 video sequences, spanning more than 560 object classes. It provides strict one-shot evaluation settings and diverse real-world tracking scenarios, making it a widely used benchmark for developing and assessing tracking models.
$\delta_{\text{iou}}$ SAM 2 0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
SA-V 76.3 80.0 80.7 80.6 80.8 80.7 81.0 80.6 80.8 80.0 77.8
LVOS 83.0 84.0 85.1 85.6 85.4 85.4 85.1 85.6 85.2 84.1 83.5
Table 9: Ablation study on IoU threshold $\delta_{\text{iou}}$.
$\delta_{\text{conf}}$ SAM 2 0 1 2 3 4 5
SA-V 76.3 80.4 80.4 80.8 80.8 79.7 78.8
LVOS 83.0 84.4 84.6 85.4 85.2 84.1 83.9